OpenStudy (anonymous):

An item is on sale for $12.00. It originally was $16.50. What is the percent decrease in its price? 38% 5% 27% 73%

OpenStudy (shane_b):

You should get:\[1-\frac{12}{16.5}*100=\% \space decrease \]

OpenStudy (anonymous):

U can work this question backwards by process of elimination...it has to be less than 50% since half of 16.50 is 8 something! so, lets try option C, 16.50 x (27%) = 4.45 then, 16.5 - 4.45 = 12! So, the correct answer is C! Enjoy!

OpenStudy (shane_b):

I guess I forgot the parentheses in there...that should have been: a few moments ago \[(1-\frac{12}{16.5})*100=\% \space decrease\]

OpenStudy (anonymous):

@Shane_B! Your method is more formal! That works too!

OpenStudy (shane_b):

Either way works...on a test sometimes trying the answers is just the fastest...sometimes not :)